OpenStudy (anonymous):

Write a quadratic function that fits the given set of points. (-2, 5), (0, -3), and (3, 0)

OpenStudy (anonymous):

@CGGURUMANJUNATH can you help?

OpenStudy (anonymous):

we'll need to make a system of equations to solve this: Y = Ax^2 + Bx + C given that you have 3 points, make three equations from the standard form of the quadratic equation written above. I'll do the first point so you see what I mean: (-2, 5) combined with Y = Ax^2 + Bx + C becomes 5 = A(-2^2) + B(-2) + C you make up the next two equations for the following two points, and we'll go from there.

OpenStudy (jdoe0001):

\({\bf (-2, 5), (0, -3), (3, 0)\qquad \qquad y = ax^2+bx+c}\\ \quad \\ \begin{array}{llll} (-2, 5)\qquad x=-2\qquad y=5\\ \quad \\ 5 = a(-2)^2+b(-2)+c\implies &\color{blue}{5=4a-2b+c}\\ \quad \\ (0, -3)\qquad x=0\qquad y=-3\\ \quad \\ -3 = a(0)^2+b(0)+c\implies -3=0+0+c\implies &\color{blue}{c=-3}\\ \quad \\ (3, 0)\qquad x=3\qquad y=0\\ \quad \\ 0 = a(3)^2+b(3)+c\implies &\color{blue}{0=9a^2+3b+c} \end{array}\)

OpenStudy (anonymous):

-3 = 0^2 + 0 + C 0 = -3^2 + -3 + C

OpenStudy (jdoe0001):

yes, just keep in mind that the last point is (3, 0), not (-3, 0)

OpenStudy (anonymous):

oh okay so 0 = 3^2 + 3 + C

OpenStudy (jdoe0001):

notice the 2nd equation, pretty much gives us "c", so you can use that in the other 2 equations and make them a system of equations of 2 variables, then solve like you'd any of those

OpenStudy (jdoe0001):

hmmm...actually... the 3rd one should be only 9a, not \(9a^2\) anyhow

OpenStudy (jdoe0001):

\(\bf 4a-2b+c=5\implies 4a-2b+(-3)=5\implies 4a-2b=8\\ \quad \\ 9a+3b+c=0\implies 9a+3b+(-3)=0\implies 9a+3b=3\)

OpenStudy (anonymous):

So the answer is 0=9a+3b+c

OpenStudy (jdoe0001):

yes \({\bf (-2, 5), (0, -3), (3, 0)\qquad \qquad y = ax^2+bx+c}\\ \quad \\ \begin{array}{llll} (-2, 5)\qquad x=-2\qquad y=5\\ \quad \\ 5 = a(-2)^2+b(-2)+c\implies &\color{blue}{5=4a-2b+c}\\ \quad \\ (0, -3)\qquad x=0\qquad y=-3\\ \quad \\ -3 = a(0)^2+b(0)+c\implies -3=0+0+c\implies &\color{blue}{c=-3}\\ \quad \\ (3, 0)\qquad x=3\qquad y=0\\ \quad \\ 0 = a(3)^2+b(3)+c\implies &\color{blue}{0=9a+3b+c} \end{array}\)

OpenStudy (jdoe0001):

then you'd just need to find "a" and "b", once you find those 3 guys, a, b, c, then plug them in in the quadratic template of \(\bf y = ax^2+bx+c\)

OpenStudy (anonymous):

Think you could help me with 5 more questions like this?

OpenStudy (jdoe0001):

if I had the time, which I don't, I'd be dashing in 5mins :/

OpenStudy (anonymous):

Yeah that's okay! thank you for all of your help!

OpenStudy (jdoe0001):

yw